# SynthVLM: High-Efficiency and High-Quality Synthetic Data for Vision Language Models

### v3

@@ Abstract.
Recently, with the rise of web images, managing and understanding large-scale image datasets has become increasingly important. Vision Large Language Models (VLLMs) have recently emerged due to their robust vision-understanding capabilities. However, training these models requires vast amounts of data, posing challenges to efficiency, effectiveness, data quality, and privacy. In this paper, we introduce SynthVLM, a novel data synthesis pipeline for VLLMs. Unlike existing methods that generate captions from images, SynthVLM employs advanced diffusion models and high-quality captions to automatically generate and select high-resolution images from captions, creating precisely aligned image-text pairs. Leveraging these pairs, we achieve state-of-the-art (SoTA) performance on various vision question answering tasks, maintaining high alignment quality and preserving advanced language abilities. Moreover, SynthVLM surpasses traditional GPT-4 Vision-based caption generation methods in performance while significantly reducing computational overhead. Crucially, our method’s reliance on purely generated data ensures the preservation of privacy, achieving SoTA performance with just 100K data points (only 18% of the official dataset size). The code is made available at https://github.com/starriver030515/SynthVLM.
$\dagger$ The first two authors have equal contributions.
* Corresponding Author
@@ 1. Introduction
Among MLLMs, VLLMs achieve competitive performance in traditional multimodal tasks such as image classification (Chen et al., 2024), image understanding (Li et al., 2023b, c), and image captioning (Bai et al., 2023). Moreover, their excellent language understanding capabilities enable strong performance in text-rich tasks, such as vision question-answering (Liu et al., 2023b, a) and image-text retrieval (Chen et al., 2024).
While most existing VLLMs focus on modifying model architecture to utilize information from multiple modalities (Li et al., 2023b, c; Bai et al., 2023; Liu et al., 2023b, a; Chen et al., 2024), data also significantly impacts the success of VLLMs. For instance, Li et al. (2023d); Wang et al. (2024a) demonstrates that higher-quality training data can enhance the performance of VLLMs. The key to ensuring high-quality data lies in the precise alignment between multimodal data, such as the alignment between captions and images. With the introduction of the DataComp (Gadre et al., 2023), increasing research efforts have focused on exploring how to achieve effective caption-image alignment. Key approaches include the use of advanced alignment techniques, the development of refined annotation guidelines, and the implementation of novel training methodologies.
@@
With the advancement of generative models, data generation strategies have increasingly been utilized to achieve data creation and alignment. For example, Nguyen et al. (2023) employed BLIP2 to generate numerous image captions, achieving SoTA results on DataComp. In the domain of VLLMs, Chen et al. (2023b) utilized GPT-4 Vision to produce highly descriptive image captions, leading to significant improvements in LLaVA. The integration of these generative models has opened new avenues for enhancing data quality and alignment, further boosting VLLM performance.
Despite the notable contributions and advances in VLLMs, current data generation and alignment strategies for VLLMs ignore generating images and face the following three key challenges:
Figure 1. We compared the synthetic image-text dataset with existing datasets. As shown in (a), the generated image can avoid content such as watermarks and advertisements. In (b), the caption mentions a joicy named plate, however in the bottom right image, joicy is actually a book. The generated images better reflect the content of the captions. Additionally, the resolution of the generated images is 1024x1024, which is higher than the existing images, is more beneficial for model training and expansion.
@@
C1. Low Data Quality. Our evaluations using metrics such as CLIPScore  (Hessel et al., 2021) and training results on VLLMs reveal that existing datasets still align the modalities sub-optimally. Web images often suffer from issues such as blurriness and watermarks. Additionally, methods employing BLIP2  (Nguyen et al., 2023) for generating captions frequently result in logical inconsistencies and unclear grammar, which mislead VLLM training and degrade their language capabilities.
C2.Poor Effectiveness. It has been demonstrated in (Du et al., 2023; He et al., 2020; Chen et al., 2023c) that low-quality data can lead to poor model performance. Consequently, VLLMs are usually trained on low-quality data, the VLLMs exhibit reduced effectiveness.
@@
C3. Low Efficiency. Methods that rely on manual captioning are labor-intensive and resource-demanding. Automated solutions like ShareGPT4v  (Chen et al., 2023b), which employ GPT-4 Vision for labeling, are expensive and difficult to scale. Moreover, current strategies often necessitate the creation of extensive datasets to enhance performance, leading to significant data redundancy.
@@
C4. Security Risks. Utilizing internet-sourced data introduces numerous security and privacy concerns  (Li et al., 2023a; Das et al., 2024). They may contain personal information or copyrighted materials, posing potential legal and ethical challenges. Moreover, the inclusion of sensitive or inappropriate content within training datasets can instigate ethical issues, thereby compromising the models’ integrity and fairness.
@@
To address these issues, we introduced a new data generation pipeline: caption-to-image synthesis. We first implemented a quality selection process for high-quality caption data. Subsequently, we employed advanced diffusion models to generate images from these captions. For quality control, we used CLIPScore as the quality metric to select high-quality image-text pairs. Our data generation method achieved higher alignment between images and captions compared to existing approaches. Utilizing 100K curated synthetic data, we achieved SoTA results on multiple benchmarks, using only 18% of the official dataset size.
@@
Figure 2. With only 100k synthetic pre-training data, our SynthVLM outperforms the LLAVA 1.5 model, which is trained on 558k data.
  • New Perspective. To the best of our knowledge, we are the first to utilize Caption-to-Image strategy to construct a high-aligned training dataset for VLLMs and achieved superior performance with just 100K data points(only 18% of the official dataset size). This sets a precedent for utilizing generated images to train large-scale VLLMs, offering a solution to the limitation of data availability, quality, and privacy.
  • New Method. We proposed a new image-text pair generation pipeline to ensure high-quality data. Additionally, we proposed a new paradigm for utilizing generated images for VLLMs training.
  • SoTA Performance. (1) High Data Quality. As shown in Figure 1, our generated data achieved higher CLIPScores, indicating better image-text alignment. Additionally, we achieved higher resolution, which is beneficial for tasks requiring high-quality images. Furthermore, our generated images avoid issues such as blurriness and watermarks.
    (2) High Efficiency and Effectiveness. As shown in Figure 2, with only 18% of the high-quality pre-training data, our method outperforms the baseline that uses 100% of the data. We not only achieve SoTA performance on vision understanding tasks but also demonstrate excellent pure text task capabilities, highlighting superior modality alignment.
    (3) Data Privacy. Using generated data avoids the need for real personal images, documents, and other sensitive information, ensuring data privacy.
@@ 2.1. Diffusion Model
While DDPMs have shown promising results, several improvements have been proposed to enhance their efficiency  (Song et al., 2021; Zhu et al., 2024) and sample quality  (Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021). The superior performance of diffusion models has been leveraged in various sub-tasks, including image generation, image translation, inpainting  (Lu et al., 2022; Su et al., 2023). Our approach leverages diffusion models to create high-quality caption-image pairs specifically for VLLM training. These generated pairs not only enrich the training data but also provide a more diverse and comprehensive dataset, enabling VLLMs to learn better representations and achieve higher performance in downstream tasks such as image captioning and visual question answering.
@@ 2.2. Vision Language Models
Recent studies  (Wang et al., 2024b; Chen et al., 2023b; Liu et al., 2023b, a; Li et al., 2023f) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al., 2023b, a) and ShareGPT4V  (Chen et al., 2023b) have shown remarkable advancements in understanding and following complex instructions through instruction tuning. Although these improvements help align the vision modality and establish a solid basis for cross-modal comprehension, they require extensive datasets for training and could potentially diminish the model’s language capabilities. In this work, we propose a novel data generation strategy that leverages high-quality caption-image pairs, enabling the model to achieve state-of-the-art (SoTA) results with minimal data. By aligning during the pre-training (PT) phase, we have further enhanced the model’s language capabilities.
@@ 2.3. Data Quality and Selection
The advent of large language models has brought about a substantial increase in the volume of training data. (Touvron et al., 2023; OpenAI, 2023b) In this scenario, the quality and quantity of data become paramount. LLMs, trained on vast amounts of data, can capture subtle nuances and complex patterns in language, excelling in various natural language processing tasks. However, the increase in data volume also brings new challenges, particularly in data management, cleaning, and annotation. (Bai et al., 2024) In this section, we mainly discuss the effectiveness of data quality and data selection.
@@
: High-quality data can significantly enhance the performance of models (meta llama, 2024). As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation. (Bai et al., 2024) Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
@@
: LLMs-based methods were commonly used in data selection. (Bai et al., 2024) For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023c) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023c) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023b) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
Given the critical role of data quality and selection in enhancing model performance, our paper focuses on leveraging advanced data selection techniques to optimize caption and image-text pair quality. By employing methods that integrate LLM data selection and image-text alignment scores, we aim to efficiently identify and utilize high-quality data for VLLMs.
@@ 2.4. Data Generation
Data has always been the key driver behind the success of LLMs. Recent advancements of LLMs largely due to the availability of large-scale, diverse, and high-quality datasets for training these models  (Lin et al., 2023). However, the scarcity of data and the high costs present substantial challenges in obtaining such datasets  (Xie et al., 2024; Xu et al., 2023a). Recent advancements in generating synthetic data and improving the performance of LLMs have shown promising results across various domains. Synthetic data holds great potential in building large-scale, high-quality datasets. Researchers have explored multiple approaches, from leveraging differential privacy to creating instruction-tuning frameworks, to enhance the quality, diversity, and utility of synthetic data  (Wang et al., 2024c; Wei et al., 2023; Lou et al., 2023; Yang et al., 2023). A key component in generating high-quality synthetic datasets is precise alignment. Fan et al.  (Fan et al., 2024) introduce REALIGN, a method that enhances the quality of instruction data by reformatting responses to better align with pre-established criteria and evidence, thereby improving LLMs’ alignment with human values while minimizing human annotation and model hallucinations. Li et al.  (Li et al., 2023e) build a high-quality instruction-following language model by automatically labeling human-written text with corresponding instructions and demonstrating highly effective self-alignment.
In the field of VLLMs, the task of constructing generative datasets has been relatively underexplored. VLLMs primarily focus on the alignment between images and captions. Existing approaches, such as ShareGPT4V  (Chen et al., 2023b), leverage GPT4-Vision to generate high-quality captions for images, thereby achieving alignment and producing SoTA results on models like LLaVA. However, this method incurs high costs and often results in suboptimal alignment due to the complexity of the captions. Our approach introduces a new method for aligning images and captions by utilizing text-to-image models to generate large-scale, high-quality data. Specifically, our method surpasses existing techniques in alignment accuracy and efficiency.
Figure 3. Curated image-text pair generation pipeline. We first use the diffusion model to generate images and then select the high clip score image-text pairs.
@@ 3. Method
In this section, we introduce our data generation pipeline and then compare the generated dataset with other commonly used datasets. In subsection 3.1, we explore the construction of our caption set and the subsequent generation of 1 million caption-image pairs using diffusion models. In subsection 3.2, we outline the filtering process applied to the generated 1 million image-text pairs. Then we meticulously select 100k high-quality, well-aligned caption-image pairs, ensuring the robustness and relevance of our dataset. In subsection 3.3, we demonstrate the effectiveness of our method by comparing the SynthVLM dataset with other existing datasets. In subsection 3.4, we summarize the VLLMs training pipeline by utilizing the synthetic dataset.
3.1. Synthetic Dataset Construction
@@
Data Source. To ensure the diversity of the captions, we combined human-generated and machine-generated captions. As shown in Table  1. The human-generated captions were primarily sourced from LAION, CC, and SBU, while the machine-generated captions were predominantly created using the method described in  (Nguyen et al., 2023). They utilize BLIP2 to regenerate captions for images in the datacomp dataset  (Gadre et al., 2023).
@@
Caption Curation. To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. This filtering process was performed using GPT-3, ensuring that only high-quality, informative captions were used for training. For the remaining captions, we calculated the CLIPScore for these captions and their corresponding raw images. CLIPScore is a metric that measures the cosine similarity between images and their corresponding captions.
The formula for calculating CLIPScore is as follows:
$\text{CLIPScore}(I,C)=\frac{\text{CLIP}(I)\cdot\text{CLIP}(C)}{||\text{CLIP}(I)||\cdot||\text{CLIP}(C)||}$
@@
where $I$ represents the image, $C$ represents the caption, and $\text{CLIP}(I)$ and $\text{CLIP}(C)$ denote the image and text feature vectors extracted by the CLIP model. The dot product of the vectors is denoted by $\cdot$, and |||||$\cdot$||| | ⋅ | | denotes the norm of the vectors.
@@
We selected the top 40% of image-caption pairs with the highest scores. These selected captions were included in the candidate caption set. At this step, we believe that captions with better image-caption alignment are more likely to generate high-quality, well-aligned images. Ultimately, we sampled a dataset of 1,000k captions for data generation. By using only captions, our method significantly reduces storage overhead and processing time. The caption curation pipeline is summarized in Figure 3(a).
@@
Image Generation. After filtering through 1000k high-quality captions, we utilized Stable Diffusion XL (SDXL)  (Podell et al., 2023), a SoTA model which can efficiently generate high-quality, high-resolution images. By configuring SDXL’s steps to 60 and utilizing 8 A100 GPUs, we were able to generate all images within a week. Furthermore, SynthVLM generates images at a resolution of 1024x1024, effectively addressing the issue of low resolution found in existing datasets. This enhancement significantly improves the quality and utility of the training data for a variety of image generation and recognition tasks.
3.2. Synthetic Data Selection
In this section, we introduce the quality control process for synthetic datasets. To further ensure the alignment between images and their corresponding text descriptions, we employ CLIPScore a second time, evaluating the quality of image-text pairs with enhanced precision.

### v4

# SynthVLM: Towards High-Quality and Efficient Synthesis of Image-Caption Datasets for Vision-Language Models

@@ Abstract.
Vision-Language Models (VLMs) have recently emerged, demonstrating remarkable vision-understanding capabilities. However, training these models requires large-scale datasets, which brings challenges related to efficiency, effectiveness, quality, and privacy of web data. In this paper, we introduce SynthVLM, a novel data synthesis and curation method for generating image-caption pairs. Unlike traditional methods, where captions are generated from images, SynthVLM utilizes advanced diffusion models and high-quality captions to automatically synthesize and select high-resolution images from text descriptions, thereby creating precisely aligned image-text pairs. To demonstrate the power of SynthVLM, we introduce SynthVLM-100K, a high-quality dataset consisting of 100,000 curated and synthesized image-caption pairs. In both model and human evaluations, SynthVLM-100K outperforms traditional real-world datasets. Leveraging this dataset, we develop a new family of multimodal large language models (MLLMs), SynthVLM-7B and SynthVLM-13B, which achieve state-of-the-art (SOTA) performance on various vision question-answering (VQA) tasks. Notably, our models outperform LLaVA across most metrics with only 18% pretrain data. Furthermore, SynthVLM-7B and SynthVLM-13B attain SOTA performance on the MMLU benchmark, demonstrating that the high-quality SynthVLM-100K dataset preserves language abilities. To facilitate future research, our dataset and the complete data generating and curating methods are open-sourced at https://github.com/starriver030515/SynthVLM.
Synthetic Images, Data Curation, Vision Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: KDD ’25: ACM KDD conference; August 3–7, 2025; Toronto, ON, Canadabooktitle: KDD ’25: ACM KDD conference, August 3–7, 2025, Toronto, ON, Canadaisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Among MLLMs, Vision Large Language Models (VLLMs) achieve competitive performance in traditional multimodal tasks such as image classification (Chen et al., 2024), image understanding (Li et al., 2023b, c), and image captioning (Bai et al., 2023). Most existing VLLMs rely on web images and primarily focus on modifying model architectures to better leverage information from multiple modalities (Li et al., 2023b, c; Bai et al., 2023; Liu et al., 2023b, a; Chen et al., 2024). However, high-quality web data is scarce, and the quality of data significantly impacts the performance of VLLMs. For example, Li et al. (2023d); Wang et al. (2024a) demonstrate that higher-quality training data can enhance the effectiveness of VLLMs. The key to ensuring high-quality data lies in the precise alignment between multimodal inputs, such as the alignment between captions and images.
@@
To address the scarcity of high-quality web data, data synthesis strategies have increasingly been employed, particularly with the advancement of generative models. For instance, Nguyen et al. (2023) utilized BLIP2 to generate numerous image captions, achieving SOTA results on DataComp. In the domain of VLLMs, Chen et al. (2023b) used GPT-4 Vision to produce highly descriptive image captions, resulting in significant improvements in LLaVA. The integration of these generative models has opened new avenues for enhancing data quality and alignment, further boosting VLLM performance.
Despite these notable advancements in VLLMs, the challenge of multimodal data persists, particularly when paired image and text are required. Although synthetic captions have been generated for images, the absence of generated images means that the issue of limited high-quality web images remains unresolved. This limitation gives rise to four key challenges that still need to be addressed:
C1. Low Data Quality.
@@
Our evaluations using metrics such as CLIPScore  (Hessel et al., 2021) and training results reveal that existing datasets still align the modalities sub-optimally. Web images often suffer from issues such as blurriness and watermarks. Additionally, methods employing BLIP2  (Nguyen et al., 2023) for generating captions frequently result in logical inconsistencies and unclear grammar, which mislead VLLM training and degrade their language capabilities.
C2.Poor Effectiveness.
Existing datasets frequently lack the necessary diversity and depth required to train VLLMs effectively. For example, many web-scraped datasets consist of low-quality or irrelevant content that fails to capture the complexity of natural language and visual information. Furthermore, the limited scope of current datasets often results in models that struggle to generalize across various contexts.
C3. Low Efficiency.
@@
Methods that rely on manual captioning are both labor-intensive and resource-intensive. Automated solutions such as ShareGPT4v (Chen et al., 2023b), which leverage GPT-4 Vision for labeling, are costly and difficult to scale. Additionally, many existing strategies require the creation of large datasets, such as LLAVA-558K, to boost performance, resulting in significant data redundancy.
C4. Security Risks.
@@
Utilizing web-sourced data introduces numerous security and privacy concerns (Li et al., 2023a; Das et al., 2024). They may contain personal information or copyrighted materials, posing potential legal and ethical challenges. Moreover, the inclusion of sensitive or inappropriate content within training datasets can instigate ethical issues, thereby compromising the models’ integrity and fairness.
@@
To address these challenges, we introduced a new data generation pipeline, SynthVLM. First, we implemented a quality selection process to filter high-quality caption data. Next, we employed advanced diffusion models to generate images based on these captions. For quality assurance, we proposed a new method for evaluating and selecting image-caption pairs by combining CLIPScore (Hessel et al., 2021) and SSIM (Wang et al., 2004), focusing on both image-text alignment and image quality. Our data generation approach achieved superior alignment between images and captions compared to existing methods. With 100K curated synthetic data, we attained SOTA results across multiple benchmarks, utilizing only 18% of the official LLaVA-558K dataset size.
@@
  • New Method. We propose a novel two-step image-caption data synthesis method called SynthVLM. Using SynthVLM, we create a high-quality image-caption dataset, SynthVLM-100K. We subsequently use SynthVLM-100K to train both the SynthVLM-7B and SynthVLM-13B models. Our method is both effective and efficient compared to the baseline method. Additionally, our ablation study demonstrates that both steps of SynthVLM are effective.
  • High Data Quality. SynthVLM-100K outperforms baseline datasets in both model and human evaluations. Specifically, SynthVLM-100K achieves superior CLIPScore and SSIMScore, indicating better image-text alignment and image quality. Moreover, it receives better ratings from human annotators, GPT-4, and Intern-VL, further confirming the dataset’s high-quality.
  • Strong Performance on VLLMs. We use only SynthVLM-100K as high-quality pre-training data, outperforms the baseline that uses LLaVA-558K. We achieve SOTA performance on vision-understanding tasks and demonstrate outstanding results on language tasks on MMLU benchmark.
  • Data Privacy. By using synthetic data, we eliminate the need for real personal images, documents, and other sensitive information, thereby ensuring data privacy.
@@ 2.1. Diffusion Model
While DDPMs have shown promising results, several improvements have been proposed to enhance their efficiency  (Song et al., 2021; Zhu et al., 2024) and sample quality  (Nichol and Dhariwal, 2021; Dhariwal and Nichol, 2021). The superior performance of diffusion models has been leveraged in various sub-tasks, including image generation, image translation, inpainting  (Lu et al., 2022; Su et al., 2023).
Figure 1. The pipeline of the SynthVLM data synthesis method is as follows: First, we filter high-quality image-caption pairs. Next, we synthesize high-quality data and subsequently filter them based on CLIPScore.
@@ 2.2. Vision Language Models
Recent studies  (Wang et al., 2024b; Chen et al., 2023b; Liu et al., 2023b, a; Li et al., 2023f) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al., 2023b, a) and ShareGPT4V  (Chen et al., 2023b) have shown remarkable advancements in understanding and following complex instructions through instruction tuning. Although these improvements help align the vision modality and establish a solid basis for cross-modal comprehension, they require extensive datasets for training and could potentially diminish the model’s language capabilities. In this work, we propose a new data generation strategy that leverages high-quality caption-image pairs, enabling the model to achieve state-of-the-art (SOTA) results with minimal data. By aligning during the pre-training (PT) phase, we have further enhanced the model’s language capabilities.
@@ 2.3. Data Quality and Selection
The advent of large language models has brought about a substantial increase in the volume of training data (Touvron et al., 2023; OpenAI, 2023b). In this scenario, the quality and quantity of data become paramount. LLMs, trained on vast amounts of data, can capture subtle nuances and complex patterns in language, excelling in various natural language processing tasks. However, the increase in data volume also brings new challenges, particularly in data management, cleaning, and annotation (Bai et al., 2024). In this section, we mainly discuss the effectiveness of data quality and data selection.
@@
: High-quality data can significantly enhance the performance of models (meta llama, 2024). As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation (Bai et al., 2024). Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
@@
: LLMs-based methods were commonly used in data selection (Bai et al., 2024). For instance, Du et al. (2023) leverages DeBERTa (He et al., 2020) for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. (2023c) score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. (2023b) use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality. Liu et al. (2023c) train two models using ChatGPT’s labeled data to score the quality and complexity of the data. Lu et al. (2023b) rely on ChatGPT to tag each instance, defining its complexity and diversity based on these tags. Parkar et al. (2024) first cluster the data, and then use GPT-4 to select high-quality data for each cluster.
Figure 2. We compared the SynthVLM-100K dataset with LLaVA-558K. As shown in (a), the generated image can avoid content such as watermarks and advertisements. In (b), the caption mentions a joicy named plate, however in the bottom right image, joicy is actually a book. The generated images better reflect the content of the captions. Additionally, the resolution of the generated images is 1024x1024, which is higher than the existing images, is more beneficial for model training and expansion.
@@ 2.4. Data Generation
Data has always been the key driver behind the success of LLMs. Recent advancements of LLMs largely due to the availability of large-scale, diverse, and high-quality datasets for training these models (Lin et al., 2023; Li et al., 2024c, e, 2023g, a, d, b). However, the scarcity of data and the high costs present substantial challenges in obtaining such datasets (Xie et al., 2024; Xu et al., 2023a). Recent advancements in generating synthetic data and improving the performance of LLMs have shown promising results across various domains. Synthetic data holds great potential in building large-scale, high-quality datasets. Researchers have explored multiple approaches, from leveraging differential privacy to creating instruction-tuning frameworks, to enhance the quality, diversity, and utility of synthetic data  (Wang et al., 2024c; Wei et al., 2023; Lou et al., 2023; Yang et al., 2023). A key component in generating high-quality synthetic datasets is precise alignment. Fan et al. (Fan et al., 2024) introduce REALIGN, a method that enhances the quality of instruction data by reformatting responses to better align with pre-established criteria and evidence, thereby improving LLMs’ alignment with human values while minimizing human annotation and model hallucinations. Li et al. (Li et al., 2023e) build a high-quality instruction-following language model by automatically labeling human-written text with corresponding instructions and demonstrating highly effective self-alignment.
@@ 3. Method
In this section, we introduce our data generation pipeline and then compare the generated dataset with other commonly used datasets. In subsection 3.1, we explore the construction of our caption set and the subsequent generation of 1M caption-image pairs using diffusion models. In subsection 3.2, we outline the filtering process applied to the generated 1 million image-text pairs. Then we meticulously select 100K high-quality, well-aligned caption-image pairs, ensuring the robustness and relevance of our dataset. In subsection 3.3, we demonstrate the effectiveness of our method by comparing the SynthVLM-100K dataset with other existing datasets.
3.1. Step1: Synthetic Dataset Construction
@@
To ensure the diversity of the captions, we combined human-generated and model-generated captions. As shown in Table  1. The human-generated captions were primarily sourced from LAION, CC, and SBU, while the model-generated captions were predominantly created using the method described in  (Nguyen et al., 2023), which utilizes BLIP2 to regenerate captions for images in the DataComp dataset  (Gadre et al., 2023).
Caption Curation.
@@
To maintain dataset quality, we first removed low-quality captions, such as advertisements, overly repetitive descriptions, and captions with significant grammatical errors. The filtering process was performed with ChatGPT, combined with statistical indicators such as N-grams and Perplexity, ensuring that only high-quality, informative captions were used for training. For the remaining captions, we calculated the CLIPScore for these captions and their corresponding raw images. CLIPScore is a metric that measures the cosine similarity between images and their corresponding captions. The formula for calculating CLIPScore is as follows:
(1) ${CLIPScore}(I,C)=\frac{{CLIP}(I)\cdot{CLIP}(C)}{||{CLIP}(I)||\cdot||{CLIP}(C)||}$
@@
where $I$ represents the image, $C$ represents the caption, and ${CLIP}(I)$ and ${CLIP}(C)$ denote the image and text feature vectors extracted by the CLIP model. The dot product of the vectors is denoted by $\cdot$, and |||||$\cdot$||| | ⋅ | | denotes the norm of the vectors.
@@
We selected the top 40% of image-caption pairs with the highest CLIPScores. These selected captions were included in the candidate caption set. Ultimately, we curated a dataset of 1M captions for data generation. By using only captions, our method significantly reduces storage overhead and processing time. The caption curation pipeline is summarized in Figure 1(a).
Image Generation.
@@
After filtering 1M high-quality captions, we employed Stable Diffusion XL (SDXL)  (Podell et al., 2023), a SOTA model capable of efficiently generating high-quality, high-resolution images. We configured SDXL with 60 sampling steps and used 8 A100 GPUs to ensure the superior quality of our synthetic images. Additionally, SynthVLM produces images at a resolution of 1024x1024, effectively addressing the low-resolution issues present in existing datasets. This improvement greatly enhances the quality and utility of the training data across various image generation and recognition tasks.
3.2. Step2: Synthetic Data Selection
In this section, we introduce a novel algorithm for quality control of generated datasets. To better ensure the alignment between images and their corresponding text descriptions, we continue to use CLIPScore. For a given image $I$ and its corresponding text $C$, we first calculate the image-text CLIPScore, ${CLIPScore}(I,C)$, to assess their alignment. Additionally, the generated images have a resolution of $1024\times 1024$, which will be resized to $336\times 336$ for compatibility with CLIP. To account for the potential loss in quality due to this resizing process, we introduce the Structural Similarity Index Measure (SSIM), a metric for image quality control. This will help us better ensure the image quality is preserved after resizing. The SSIM formula is defined as follows:
$SSIM(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$
Where $x$ and $y$ are the two images being compared, $\mu_{x}$ and $\mu_{y}$ are the mean pixel intensities of the images, $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are the variances of the images, $\sigma_{xy}$ is the covariance of the images, $C_{1}$ and $C_{2}$ are constants to stabilize the division with weak denominator values.
For a given image $I$ with a resolution of $1024\times 1024$, we first resize it to $336\times 336$, then interpolate to restore it back to $1024\times 1024$. The SSIM value is then computed between the resized image and the original image to quantify the loss introduced by the resizing process. Let $I_{{resized}}$ denote the resized image, where:
$I_{{resized}}={resize}({resize}(I,(336,336)),(1024,1024))$
The SSIMScore is computed as:
$SSIMScore=SSIM(I,I_{{resized}})$
Finally, to combine the CLIPScore and SSIMScore, we apply a weighted sum, with SSIMScore weighted by a factor $\lambda$. The overall score $S$ is given by:
$Weighted\_Score={CLIPScore}(I,C)+\lambda\cdot SSIMScore(I,I_{{resized}})$
In practice, we set $\lambda=0.5$ to balance the contribution of CLIPScore and SSIMScore in the final score.